A NIST standard solution of mercury(II) chloride (HgCl2) in water has the true concentration of   [Hg2+] = 4.50 ppm. Researcher A measured [Hg2+] for this solution in triplicate, and obtained the following values: 4.20 ppm, 4.40 ppm, 4.65 ppm. Researcher B also measured [Hg2+] in triplicate, and obtained the following values: 3.95 ppm, 4.00 ppm, 4.02 ppm. Which researcher’s measurements were more precise? Which researcher’s measurements were more accurate? Explain your answer with calculations.
Answer #1

Descriptive Statistics: reasercher A

Variable N Mean SE Mean StDev Minimum Q1 Median Q3 Maximum
reseacher A 3 3.9900 0.0208 0.0361 3.9500 3.9500 4.0000 4.0200 4.0200


Descriptive Statistics: researcher B

Variable N Mean SE Mean StDev Minimum Q1 Median Q3 Maximum
reseacher B 3 4.417 0.130 0.225 4.200 4.200 4.400 4.650 4.650

since we know that Precision is inversely related to standard error. When the standard error is small, sample estimates are more precise. hence researcher A 's measurements were more precise.

now Accuracy is the how close you are to the true value.

range of researcher A 4.0200 -3.9500 = 0.07

range of researcher B = 4.650 - 4.200 = 0.45

researcher A  has the lowest range (.07) and so is the more precise.